OpenStudy (anonymous):

4x^2-7x+1 whats the number of real solutions? and whats the discriminant

OpenStudy (anonymous):

zero real solutions... the discriminant is -2

OpenStudy (anonymous):

the dis is 33 i think

OpenStudy (kinggeorge):

Since the discriminant is \((-7)^2-4(4)(1) = 49-16=33\), and the greatest exponent is 2, it must have 2 real solutions.

OpenStudy (anonymous):

No... the equation for the discriminant is: \[b^2-4ac\]

OpenStudy (anonymous):

ohhh.. right.. wow...

OpenStudy (anonymous):

anyways.. you have to formula for the discriminant hahah :)

OpenStudy (anonymous):

there are two real solutions is the number is greater than zero.. if the discriminant is zero then there is one solution.. if it's less than zero then there are no real solutions.

OpenStudy (kinggeorge):

A positive discriminant means 2 real root, a negative discriminant means 0 real roots, and a discriminant equal to 0 means 1 real root. What madscientist said...

OpenStudy (anonymous):

tx

OpenStudy (kinggeorge):

you're welcome.